OpenStudy (anonymous):

Let F(x)=ln(1+x^2) A) limit as x approaches infinity B)F1(x) as it approaches infinity C) explain the behavior as function x gets large

OpenStudy (zzr0ck3r):

what is \(\ln(1+\infty^2)\)?

OpenStudy (anonymous):

you mean the derivative?

OpenStudy (zzr0ck3r):

no we are just doing limits, not derivatives. Part asks you what happens to \(\ln(1+x^2)\) when x gets very large. So what happens when you "plug in" infinity?

OpenStudy (zzr0ck3r):

Part A asks*

OpenStudy (anonymous):

the number becomes particularly small ?

OpenStudy (anonymous):

for example ln (1+1000^2)=13.8155

OpenStudy (zzr0ck3r):

make it 10^50

OpenStudy (zzr0ck3r):

its actually getting larger, so it will tend to infinity. VERY SLOWLY......

OpenStudy (anonymous):

115.129

OpenStudy (zzr0ck3r):

no try 10 ^1000000000000000000000

OpenStudy (zzr0ck3r):

now try*

OpenStudy (anonymous):

my calculator wont allow me but i guess its bc it is very large

OpenStudy (zzr0ck3r):

yes, it keeps getting larger

OpenStudy (zzr0ck3r):

what is F1(x)? Do you mean \(f^{-1}(x)\)?

OpenStudy (anonymous):

no first derivative as it approaches infinity

OpenStudy (zzr0ck3r):

ahh good, otherwise we were in trouble.

OpenStudy (zzr0ck3r):

do you know the first derivative?

OpenStudy (anonymous):

so for part A i would put infinity ?

OpenStudy (zzr0ck3r):

yes, or undefined

OpenStudy (anonymous):

2x/x^2+1

OpenStudy (anonymous):

that is the derivative

OpenStudy (zzr0ck3r):

sorry, phone. ok do you know how to take the limit?

OpenStudy (anonymous):

it confuses me a little ?

OpenStudy (anonymous):

my teacher alwasy taught us to just plug it in

OpenStudy (zzr0ck3r):

well you have \(\frac{2x}{x^2+1}\) if you divide every term by x you get \(\frac{2}{x+\frac{1}{x}}\) if you take the limit as x goes to infinity you get \(\frac{2}{\infty+\frac{1}{\infty}}=\frac{2}{\infty+0}=\frac{2}{\infty}=?\)

OpenStudy (anonymous):

a very small number

OpenStudy (anonymous):

approaches zero ?